A steel tank filled with chlorine gas has an internal gas pressure of 13.6 bar at 24.0oC.
Calculate the temperature to which this gas sample must be heated for it to register a pressure of 1.57 x 103 kPa
1
Expert's answer
2020-04-06T11:50:56-0400

PV = nRT

n = PV / RT

V1 = V2, R = const,

P1 / T1 = P2 / T2

T2 = P2T1 / P1

T2 = 1.57 × 106 × (24+273.15) / 1.36 × 106 = 343 oK = 70 oC

Answer: 70 oC.